# NeurOp-Diff: Continuous Remote Sensing Image Super-Resolution via Neural Operator Diffusion

### v2

@@ Abstract
Most publicly accessible remote sensing data suffer from low resolution, limiting their practical applications. To address this, we propose a diffusion model guided by neural operators for continuous remote sensing image super-resolution (NeurOp-Diff). Neural operators are used to learn resolution representations at arbitrary scales, encoding low-resolution (LR) images into high-dimensional features, which are then used as prior conditions to guide the diffusion model for denoising. This effectively addresses the artifacts and excessive smoothing issues present in existing super-resolution (SR) methods, enabling the generation of high-quality, continuous super-resolution images. Specifically, we adjust the super-resolution scale by a scaling factor $s$, allowing the model to adapt to different super-resolution magnifications. Furthermore, experiments on multiple datasets demonstrate the effectiveness of NeurOp-Diff. Our code is available at https://github.com/zerono000/NeurOp-Diff.
@@ 1 Introduction
Remote sensing super-resolution (RSSR) is a key technology in remote sensing image processing [51], aiming to restore low-resolution remote sensing (LRRS) images to high resolution remote sensing (HRRS) images, thus enhancing spatial details in remote sensing imagery [39]. Despite advances in remote sensing technology that have greatly improved the spatial resolution of modern satellites, challenges related to resolution limitations persist [2]. On the other hand, different application scenarios, such as urban planning and agricultural monitoring, have varying requirements for image details and coverage. Continuous super-resolution can dynamically adjust the magnification factor according to specific needs [6, 42], providing greater adaptability. Moreover, remote sensing data comes from diverse sources with varying resolutions. Continuous super-resolution helps integrate data of different resolutions, improving the efficiency and quality of data fusion.
@@
Early regression-based methods performed image super-resolution by establishing a mapping between low-resolution (LR) and high-resolution (HR) images. These approaches often relied on deep convolutional neural networks (CNNs) [8, 9, 22, 29, 37, 48]. The advantages of these methods include stable training, simple and easy-to-optimize loss functions, typically using mean squared error (MSE) to minimize pixel-wise differences between the reconstructed and ground-truth images. However, regression-based methods often result in overly smooth outputs that lack detail and struggle to recover realistic high-frequency information. On the other hand, variational autoencoders (VAEs) [21, 24] and generative adversarial networks (GANs) [12, 19, 20, 26, 32, 40] have demonstrated exceptional image generation capabilities. GANs, in particular, introduce an adversarial learning mechanism to produce more realistic high-resolution images. Compared to regression-based methods, GAN-based approaches better preserve image details, generating sharper and more natural results. However, the adversarial training process is unstable, requiring careful hyperparameter tuning and often encountering mode collapse issues. Moreover, GAN-generated images may suffer from artifacts, leading to distortions in specific areas of the reconstructed images [12].
@@
Recently, diffusion models (DMs) [15, 35] have gained attention for their strong performance in generating high-quality images. Compared to GANs, diffusion models exhibit greater stability during training and can generate more complex and realistic image structures. In the context of image super-resolution, methods like SR3 iteratively refine high-resolution images by progressively denoising from noisy inputs [34], achieving excellent reconstruction results. However, such methods are constrained by fixed scaling factors [7, 16, 50], requiring separate models for different magnification scales, which increases computational cost. Furthermore, due to the limited integration of additional prior knowledge, their performance still has room for improvement.
@@
In this paper, we propose the diffusion model using neural operators as conditional guidance(NeurOp-Diff), for continuous super-resolution of remote sensing images. By leveraging the neural operator’s ability to learn mappings between infinite-dimensional function spaces, we achieve continuous super-resolution for remote sensing images. The LRRS images processed by the neural operator serve as additional priors for the model’s denoising process. Unlike previous works that used bicubic upsampled LR images as prior knowledge, the neural operator dynamically updates latent bases through a multilayer Galerkin-type attention mechanism [42]. This process helps to recover certain high-frequency details, providing priors that enable the generation of more realistic high-resolution images [33] (see Figure 1). Experimental results demonstrate that the images generated by NeurOp-Diff surpass those produced by previous methods, both in quality and detail.
@@
  • We propose a joint framework that integrates neural operators with diffusion models and, for the first time, applies neural operators to the task of remote sensing image super-resolution.
@@
  • We conducted extensive experiments on three remote sensing datasets (i.e., UCMerced, AID and RSSCN7), and compared with six advanced methods (e.g., SR3, IDM, EDiffSR, etc.), our results showed greater competitiveness.
@@ 2 Relate Work
Remote Sensing Super-Resolution Generation Models. Deep convolutional neural networks (CNNs) [8, 9, 22, 29, 37, 48] were among the earliest and most prominent methods for natural image super-resolution and have since been adapted to remote sensing image super-resolution(SR) due to their impressive performance. For example, LGCNet [43] uses a ”multi-branch” structure to learn both local details and global environmental priors of remote sensing images, while RDBPN [49]employs a residual dense back-projection network to achieve large-scale super-resolution reconstruction.With the successful application of Transformers in computer vision [5, 10, 13, 30, 38, 41], Transformer-based super-resolution methods have started to gain significant attention. Although the SwinIR [28] model has shown excellent image reconstruction performance, it tends to struggle when it comes to restoring high-frequency details and complex textures. Simultaneously, Generative Adversarial Networks(GANs) [12, 19, 20, 26, 32, 40] have shown strong potential for restoring image details. For example, CDGAN [1] enhances model discrimination by introducing a dual-path network architecture, stochastic gates, and coupled adversarial loss, while EEGAN [18] employs a dense edge-enhancement network to reduce noise and improve edge contours. However, GANs are known for their unstable training process, which can lead to model collapse.
@@
In recent years, denoising diffusion models(DDMs) [15, 35, 36] have emerged as a leading approach in the image generation domain. The denoising diffusion probabilistic model (DDPM) [15] introduced a new paradigm for image generation. Subsequent models such as Stable Diffusion [33] showcased the powerful capabilities of diffusion models in text-to-image generation, garnering significant attention. In specific tasks like image super-resolution, generative models have demonstrated unique advantages. For example, SR3 [34] achieves impressive super-resolution results by iterative refinement. However, its performance remains limited by insufficient prior knowledge. To address this, we extend SR3 by introducing neural operators, which enable continuous super-resolution while providing richer prior information. This enhancement significantly improves the denoising capability of the diffusion model and overall image quality.
@@
Neural Operators. Neural operators (NO) are an emerging deep learning model capable of effectively learning and approximating complex nonlinear mappings while preserving the physical invariance and geometric structure of problems. This makes them particularly well-suited for solving complex partial differential equations (PDEs) [14, 17, 27, 31]. Unlike traditional neural networks, which typically learn mappings between finite-dimensional spaces, neural operators are designed to learn mappings from function space to function space, unconstrained by spatial dimensions. They can handle inputs of varying resolutions and sizes, producing corresponding functional outputs.
@@
In recent years, neural operators have been applied to super-resolution tasks with promising results. For example, SRNO [42] employs a Galerkin-type attention mechanism and a multilayer architecture to treat the mapping from low-resolution to high-resolution images as a transformation in continuous function space, enabling fast, efficient, and dynamically adaptive super-resolution reconstruction. However, such methods often exhibit over-smoothing when restoring high-frequency details and textures in images.To address this, we integrate denoising diffusion models into our work. Diffusion models are capable of generating high-quality images with rich details and natural textures. By combining diffusion models with neural operators, our approach retains the advantage of neural operators in achieving continuous super-resolution while mitigating the issue of over-smoothing in image reconstruction.
@@ 3 Method
Given a set of remote sensing image datasets $D=\{x_{i},y_{i},z_{i}\}^{N}_{i=1}$ and a control output resolution factor $s$, the datasets satisfy a specific joint probability distribution $p(X,Y,Z)$. $x_{i}$ represents low-resolution source images, and $y_{i}$ represents high-resolution degraded images obtained through the network. Using the approximation of the parameters $p(z|y)$, we aim to map low-resolution images to high-resolution images.
NeurOp-Diff first initializes $z_{i}$ as pure noise $z_{T}\sim\mathcal{N}(0,I)$, and then refines it iteratively $T$ times through the learned conditional transition distribution $p_{\theta}(z_{t-1}\mid z_{t},y)$ to generate high-resolution images $z_{0}\sim p(z\mid y)$. The forward process defines how to progressively add noise to the initial $z_{0}$, generating the intermediate state sequence $(z_{T},z_{T-1},\ldots,z_{0})$ via a fixed Markov transition probability $q(z_{t}\mid z_{t-1})$. Our goal is to reverse this diffusion process using the reverse chain based on $y$, achieving the task of recovering the signal from noisy data. The reverse chain employs a denoising model built upon a U-Net architecture, taking as input low-resolution images processed by neural operators and noisy target images. Details of the model are as follows (see Figure 2).
@@ 3.1.1 Forward Process
We first define a forward Markov diffusion process $q$ [15], which progressively adds noise to the initial image $z_{0}$:
@@
where $\alpha_{t}$ is a hyperparameter controlling the noise level, satisfying $10<\alpha_{t}<$ 1. By marginalizing over intermediate steps, we can derive the boundary distribution of $z_{t}$ given $z_{0}$:
@@
where $\gamma_{t}=\prod_{i=1}^{t}\alpha_{i}$. Using the properties of Gaussian distributions, the posterior distribution $q(z_{t-1}\mid z_{0},z_{t})$ can be derived as:
@@ 3.1.2 Reverse Process
By learning the conditional distribution of the forward diffusion process $p_{\theta}(z_{t-1}\mid z_{t},y)$, the inference process can approximate the true posterior distribution [35], the final optimization step of the model is:
@@ 3.2 Neural Operator
Neural operators map input functions to their latent representations and use kernel integration to capture global correlations in the function space. Specifically, we define a kernel function $K_{t}:\mathbb{R}^{d_{r}+d_{r}}\to\mathbb{R}^{d_{r}\times d_{r}}$, which operates on the hidden representations $\phi(x^{\prime})$ and $\phi(y^{\prime})$, rather than relying on spatial variables $(x^{\prime},y^{\prime})$. To efficiently explore the hidden representations $\phi(x^{\prime})$, we can use multiple test functions to define these distributions, similar to single-head self-attention [25].
The kernel integral operator can be approximated using Monte Carlo methods, as follows:
[…]
where the kernel function $K(\phi(x^{\prime}),\phi_{i})$ is computed as:
$K(\phi(x^{\prime}),\phi_{i})=\frac{\exp\left(\frac{\langle W_{q}\phi(x^{\prime}),W_{k}\phi_{i}\rangle}{\sqrt{d_{r}}}\right)}{\sum_{j=1}^{n_{f}}\exp\left(\frac{\langle W_{q}\phi_{j},W_{k}\phi_{i}\rangle}{\sqrt{d_{r}}}\right)}W_{v}$ (9)
@@
The advantages of neural operators lie in their ability to handle data with different resolutions while capturing global relationships in the function space through inner product operations. For example, a function $f$ can be represented as a distribution function instead of being evaluated at individual coordinate points. We use query, key, and value functions $q(x^{\prime})$, $k(x^{\prime})$, and $v(x^{\prime})$ to express global correlations, and the kernel integration operation can be represented as:
$((\mathcal{K}(\phi))(x^{\prime}))_{j}=\sum_{l=1}^{d_{r}}\langle k_{l},v_{j}\rangle q_{l}(x^{\prime}).$ (10)
In the framework of neural operators, the latent representation $\phi$ is obtained through a linear combination of basis functions. To achieve efficient computation in high-resolution scenarios, Galerkin-type attention [4] is introduced. This mechanism achieves an approximately optimal solution through the linear combination of $Q$, $K$, and $V$:
@@
where $\tilde{K}$ and $\tilde{V}$ are normalized keys and values. This method not only reduces computational complexity but also effectively addresses discontinuities at feature boundaries. To further enhance expressive power, neural operators iteratively update representations through a feed-forward network (FFN). The iterative process is described as:
$\phi_{t+1}(x^{\prime})=\phi_{t}(x^{\prime})+\mathcal{O}\left((K_{t}(\phi_{t}))(x^{\prime})+\phi_{t}(x^{\prime})\right).$ (12)
This iterative mechanism enables neural operators to continuously enrich their representational capabilities, improving performance in ultra-high-resolution tasks.

### v3

@@ Abstract
Most publicly accessible remote sensing data suffer from low resolution, limiting their practical applications. To address this, we propose a diffusion model guided by neural operators (NO) for continuous remote sensing image super-resolution (NeurOp-Diff). Neural operators are used to learn resolution representations at arbitrary scales, encoding low-resolution (LR) images into high-dimensional features, which are then used as prior conditions to guide the diffusion model for denoising. This effectively addresses the artifacts and excessive smoothing issues present in existing super-resolution (SR) methods, enabling the generation of high-quality, continuous super-resolution images. Specifically, we adjust the super-resolution scale by a scaling factor $s$, allowing the model to adapt to different super-resolution magnifications. Furthermore, experiments on multiple datasets demonstrate the effectiveness of NeurOp-Diff. Our code is available at https://github.com/zerono000/NeurOp-Diff.
@@ 1 Introduction
Remote sensing super-resolution (RSSR) is a key technology in remote sensing image processing [zhu2017deep], aiming to restore low-resolution remote sensing (LRRS) images to high resolution remote sensing (HRRS) images, thus enhancing spatial details in remote sensing imagery [wang2022comprehensive]. Despite advances in remote sensing technology that have greatly improved the spatial resolution of modern satellites, challenges related to resolution limitations persist [bashir2021comprehensive]. On the other hand, different application scenarios, such as urban planning and agricultural monitoring, have varying requirements for image details and coverage. Continuous super-resolution can dynamically adjust the magnification factor according to specific needs [chen2021learning, wei2023super], providing greater adaptability. Moreover, remote sensing data comes from diverse sources with varying resolutions. Continuous super-resolution helps integrate data of different resolutions, improving the efficiency and quality of data fusion.
@@
Early regression-based methods performed image super-resolution by establishing a mapping between low-resolution (LR) and high-resolution (HR) images. These approaches often relied on deep convolutional neural networks (CNNs) [Dong2015, Dong2016, Kim2016a, Lim2017, vahdat2020nvae, Zhang2018]. The advantages of these methods include stable training, simple and easy-to-optimize loss functions, typically using mean squared error (MSE) to minimize pixel-wise differences between the reconstructed and ground-truth images. However, regression-based methods often result in overly smooth outputs that lack detail and struggle to recover realistic high-frequency information. On the other hand, variational autoencoders (VAEs) [Kim2016b, Kingma2013] and generative adversarial networks (GANs) [goodfellow2014generative, Karras2018, karras2019style, ledig2017photo, wang2018esrgan, dong2021rrsgan] have demonstrated exceptional image generation capabilities. GANs, in particular, introduce an adversarial learning mechanism to produce more realistic high-resolution images. Compared to regression-based methods, GANs-based approaches better preserve image details, generating sharper and more natural results. However, the adversarial training process is unstable, requiring careful hyperparameter tuning and often encountering mode collapse issues. Moreover, GANs generated images may suffer from artifacts, leading to distortions in specific areas of the reconstructed images [goodfellow2014generative].
@@
Recently, diffusion models (DMs) [ho2020denoising, sohl2015deep] have gained attention for their strong performance in generating high-quality images. Compared to GANs, diffusion models exhibit greater stability during training and can generate more complex and realistic image structures. In the context of image super-resolution, methods like SR3 [saharia2022image] iteratively refine high-resolution images by progressively denoising from noisy inputs, achieving excellent reconstruction results. However, such methods are constrained by fixed scaling factors [dai2019second, hsu2024drct, zhou2023learning], requiring separate models for different magnification scales, which increases computational cost. Furthermore, due to the limited integration of additional prior knowledge, their performance still has room for improvement.
@@
In this paper, we propose the diffusion model using neural operators as conditional guidance (NeurOp-Diff), for continuous super-resolution of remote sensing images. By leveraging the neural operator’s ability to learn mappings between infinite-dimensional function spaces, we achieve continuous super-resolution for remote sensing images. The LRRS images processed by the neural operator serve as additional priors for the model’s denoising process. Unlike previous works that used bicubic upsampled LR images as prior knowledge, the neural operator dynamically updates latent bases through a multilayer Galerkin-type attention mechanism [wei2023super]. This process helps to recover certain high-frequency details, providing priors that enable the generation of more realistic high-resolution images [rombach2022high] (see Figure 1). Experimental results demonstrate that the images generated by NeurOp-Diff surpass those produced by previous methods, both in quality and detail.
@@
  • We propose a unified frame-work that integrates neural operators with diffusion models to remote sensing image super-resolution for the first time.
@@
  • We conducted extensive experiments on three remote sensing datasets, and compared with five advanced methods, our results showed greater competitiveness.
@@ 2 Relate Work
Remote Sensing Super-Resolution Generation Models. Deep convolutional neural networks (CNNs) [Dong2015, Dong2016, Kim2016a, Lim2017, vahdat2020nvae, Zhang2018] were among the earliest and most prominent methods for natural image super-resolution and have since been adapted to remote sensing image super-resolution (SR) due to their impressive performance. For example, LGCNet [wu2023] uses a multi-branch structure to learn both local details and global environmental priors of remote sensing images, while RDBPN [zhang2018residual] employs a residual dense back-projection network to achieve large-scale super-resolution reconstruction. With the successful application of Transformers in computer vision [chen2021pre, dosovitskiy2020image, guo2022image, liu2021swin, vaswani2017attention, wang2022uformer], Transformer-based super-resolution methods have started to gain significant attention. Although the SwinIR [liang2021swinir] model has shown excellent image reconstruction performance, it tends to struggle when it comes to restoring high-frequency details and complex textures. Simultaneously, GANs [goodfellow2014generative, Karras2018, karras2019style, ledig2017photo, dong2021rrsgan, wang2018esrgan] have shown strong potential for restoring image details. For example, EEGAN [Ji2019] employs a dense edge-enhancement network to reduce noise and improve edge contours. However, GANs are known for their unstable training process, which can lead to model collapse.
@@
In recent years, denoising diffusion models (DDMs) [ho2020denoising, sohl2015deep, song2020denoising] have emerged as a leading approach in the image generation domain. The denoising diffusion probabilistic model (DDPM) [ho2020denoising] introduced a new paradigm for image generation. Subsequent models such as Stable Diffusion [rombach2022high] showcased the powerful capabilities of diffusion models in text-to-image generation, garnering significant attention. In specific tasks like image super-resolution, generative models have demonstrated unique advantages. For example, SR3 [saharia2022image] achieves impressive super-resolution results by iterative refinement. However, its performance remains limited by insufficient prior knowledge. Recent works like Ref-Diff [dong2024building] and SGDM [wang2025semantic] incorporate semantic priors—land cover changes and vector maps—into conditional diffusion or generative models, enhancing fidelity and realism in large-scale remote sensing SR. However, they are still limited to fixed upscaling factors. To address this, we introduce neural operators based on SR3, enabling not only continuous super-resolution but also providing richer prior knowledge to enhance the denoising performance of the diffusion model.
@@
Neural Operators. Neural operators are an emerging deep learning model capable of effectively learning and approximating complex nonlinear mappings while preserving the physical invariance and geometric structure of problems. This makes them particularly well-suited for solving complex partial differential equations [operator1, operator2, operator3, operator4]. Unlike traditional neural networks, which typically learn mappings between finite-dimensional spaces, neural operators are designed to learn mappings from function space to function space, unconstrained by spatial dimensions. They can handle inputs of varying resolutions and sizes, producing corresponding functional outputs.
@@
In recent years, neural operators have been applied to super-resolution tasks with promising results. For example, SRNO [wei2023super] employs a Galerkin-type attention mechanism and a multilayer architecture to treat the mapping from low-resolution to high-resolution images as a transformation in continuous function space, enabling fast, efficient, and dynamically adaptive super-resolution reconstruction. However, such methods often exhibit over-smoothing when restoring high-frequency details and textures in images. To address this, we integrate denoising diffusion models into our work. Diffusion models are capable of generating high-quality images with rich details and natural textures. By combining diffusion models with neural operators, our approach retains the advantage of neural operators in achieving continuous super-resolution while mitigating the issue of over-smoothing in image reconstruction.
@@ 3 Method
Given a set of remote sensing image datasets $D=\{(x^{(i)},y^{(i)},z^{(i)})\}_{i=1}^{N}$ and a control output resolution factor $s$, the datasets satisfy a specific joint probability distribution $p(X,Y,Z)$, where $x^{(i)}$ denotes the low-resolution source image, $y^{(i)}$ denotes the degraded high-resolution image obtained through the network, and $z^{(i)}$ denotes the ground-truth high-resolution image. Using the approximation of the conditional distribution $p(z|y)$ [wu2023hsr], we aim to map low-resolution images to their corresponding high-resolution counterparts.
NeurOp-Diff first initializes $z^{(i)}$ as pure noise $z_{T}^{(i)}\sim\mathcal{N}(0,I)$, and then refines it iteratively $T$ times through the learned conditional transition distribution $p_{\theta}(z_{t-1}^{(i)}\mid z_{t}^{(i)},y^{(i)})$ to generate high-resolution images $z_{0}^{(i)}\sim p(z\mid y)$. Here, $\theta$ denotes the learnable parameters of the neural network used to model the denoising process. The forward process defines how to progressively add noise to the initial $z_{0}^{(i)}$, generating the intermediate state sequence $(z_{T}^{(i)},z_{T-1}^{(i)},\ldots,z_{0}^{(i)})$ via a fixed Markov transition probability $q(z_{t}\mid z_{t-1})$. Our goal is to reverse this diffusion process using the reverse chain based on $y^{(i)}$, achieving the task of recovering the signal from noisy data. The reverse chain employs a denoising model built upon a U-Net architecture, taking as input low-resolution images processed by neural operators and noisy target images. Details of the model are as follows (see Figure 2).
@@ 3.1.1 Forward Process
We first define a forward Markov diffusion process [ho2020denoising] $q$, which progressively adds noise to the initial image $z_{0}$:
@@
where $\alpha_{t}$ is a hyperparameter controlling the noise level, satisfying $0<\alpha_{t}<1$. By marginalizing over intermediate steps, we can derive the boundary distribution of $z_{t}$ given $z_{0}$:
@@
where $\gamma_{t}=\prod_{i=1}^{t}\alpha_{i}$. Using the properties of Gaussian distributions, the posterior distribution $q(z_{t-1}\mid z_{0},z_{t})$ can be derived as [saharia2022image]:
@@ 3.1.2 Reverse Process
By learning the conditional distribution of the forward diffusion process $p_{\theta}(z_{t-1}\mid z_{t},y)$, the inference process can approximate the true posterior distribution [sohl2015deep], the final optimization step of the model is [saharia2022image]:
@@ 3.2 Neural Operator
Neural operators map input functions to their latent representations and use kernel integration to capture global correlations in the function space. Specifically, we define a kernel function $K_{t}:\mathbb{R}^{d_{r}+d_{r}}\to\mathbb{R}^{d_{r}\times d_{r}}$, which operates on the hidden representations $\phi(\xi)$ and $\phi(\eta)$, rather than relying on spatial variables $(\xi,\eta)$ [wei2023super]. Here, $d_{r}$ denotes the feature dimension, and $K_{t}$ represents the kernel generation function at the $t$-th layer. To efficiently explore the hidden representations $\phi(\xi)$, we will use multiple test functions to define these distributions, similar to single-head self-attention [Kovachki2021]. Let $\phi_{i}=\phi(\xi_{i})\in\mathbb{R}^{d_{r}}$, for $i=1,\ldots,n_{h_{f}}$, where $n_{h_{f}}$ denotes the number of pixels on the high-resolution grid.
The kernel integral operator $\mathcal{K}(\phi)$ can be approximated using Monte Carlo methods, as follows:
$\begin{gathered}\mathcal{K}(\phi)(\xi)=\int_{\Omega}K(\phi(\xi),\phi(\eta))\phi(\eta)\,d\eta\\ \approx\sum_{i=1}^{n_{f}}K(\phi(\xi),\phi_{i})\phi_{i},\quad\forall\xi\in\Omega_{h_{f}}\end{gathered}$ (8)
where the kernel function $K(\phi(\xi),\phi_{i})$ is computed as:
$K(\phi(\xi),\phi_{i})=\frac{\exp\left(\frac{\langle W_{q}\phi(\xi),W_{k}\phi_{i}\rangle}{\sqrt{d_{r}}}\right)}{\sum_{j=1}^{n_{f}}\exp\left(\frac{\langle W_{q}\phi_{j},W_{k}\phi_{i}\rangle}{\sqrt{d_{r}}}\right)}W_{v}$ (9)
@@
The advantages of neural operators lie in their ability to handle data with different resolutions while capturing global relationships in the function space through inner product operations [wei2023super]. For example, a function $f$ can be represented as a distribution function instead of being evaluated at individual coordinate points. We use query, key, and value functions $q(\xi)$, $k(\xi)$, and $v(\xi)$ to express global correlations, and the kernel integration operation can be represented as $(j=1,\ldots,d_{r})$:
$((\mathcal{K}(\phi))(\xi))_{j}=\sum_{l=1}^{d_{r}}\langle k_{l},v_{j}\rangle q_{l}(\xi),\quad\forall\xi\in\Omega_{h_{f}}$ (10)
In the framework of neural operators, the latent representation $\phi$ is obtained through a linear combination of basis functions. To achieve efficient computation in high-resolution scenarios, Galerkin-type attention [cao2021choose] is introduced. This mechanism achieves an approximately optimal solution through the linear combination of $Q$, $K$, and $V$:
@@
where $\tilde{K}$ and $\tilde{V}$ are normalized keys and values. This method not only reduces computational complexity but also effectively addresses discontinuities at feature boundaries. To further enhance expressive power, neural operators iteratively update representations through a feed-forward network (FFN $\mathcal{O}:\mathbb{R}^{d_{r}}\to\mathbb{R}^{d_{r}}$). The iterative process is described as:
$\phi_{t+1}(\xi)=\phi_{t}(\xi)+\mathcal{O}\left((K_{t}(\phi_{t}))(\xi)+\phi_{t}(\xi)\right).$ (12)
This iterative mechanism [wei2023super] enables neural operators to continuously enrich their representational capabilities, improving performance in ultra-high-resolution tasks.